OpenStudy (anonymous):

what is a equation that relates universal gravitational constant to the accleration of free fall at the surface of the earth

OpenStudy (egenriether):

newtons gravitational equation gives the force of attraction of two masses. One could be the earth. Then Newton's second law of motion tells what that force does to objects.

OpenStudy (anonymous):

\[T^4=(4piL/g^2) (k-2d)\] change to the form y=mx+c where y=T^4, x=d, L=180

OpenStudy (egenriether):

What? Is this a general question? Free fall at the surface of the earth usually implies neglecting the variation in gravity. http://en.wikipedia.org/wiki/Newton's_law_of_universal_gravitation